10. A single slit 1.0 mm wide is illuminated by 450 nm light. What is the central maximum (in cm) width in the diffraction pattern on a screen 5.0 m away?




A. 3 cm



B. 0.8 cm



C. 0.45 cm



D. 0.5 cm




1
Expert's answer
2021-12-30T15:20:59-0500

A single slit 1.0 mm wide is illuminated by 450 nm light. What is the central maximum (in cm) width in the diffraction pattern on a screen 5.0 m away?



"(m+1/2)\lambda=d\sin\theta,\\\space\\
\theta=0.0129°."

The width:


"x=2L\tan\theta=2.25\text{ mm}."

Two such fringes give 0.45 cm.